40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 13 19:36:36 2014, Koji, Summary, LSC, PRM 2nd/4th violin filter added 
    Reply  Tue Feb 25 18:16:22 2014, Jenne, Summary, LSC, Even more violin filters PRMvio2.png
    Reply  Tue Apr 1 21:45:01 2014, Jenne, Summary, LSC, PRM violin notch not at correct freq? 
Message ID: 9634     Entry time: Thu Feb 13 19:36:36 2014     Reply to this: 9674   9772
Author: Koji 
Type: Summary 
Category: LSC 
Subject: PRM 2nd/4th violin filter added 

Jenne and I noticed high pitch sound from our acoustic interferometer noise diagnostic system.
The frequency of this narrow band noise was 1256Hz, which is enough close to twice of the PRM violin mode freq.
After putting notch filter at 1256+/-25Hz at the violin filters, the noise is gone. Just in case I copied the same filters to all of the test masses.

Later, I found that the 4th violin modes are excited. Additional notch filters were added to "vio3" filter bank to mitigate the oscillation.

ELOG V3.1.3-